# Generalization Bounds and Model Complexity for Kolmogorov–Arnold Networks

### v1

@@ Abstract
Keywords: Empirical process, Deep neural networks, Generalization bounds, Kolmogorov-Arnold representation theorem, Model complexity
@@ 1.1 Contributions
  • We empirically investigate these bounds when the KANs are trained with stochastic gradient descent (SGD) on simulated and real data sets. The results shed new light on how the performance of KANs is affected by the underlying network structure and the network complexity. In particular, we observe that the complexity measure derived in Section 2 is tightly correlated with the excess loss, demonstrating the complexity measure’s practical relevance.
@@ 1.2 Related works
Generalization bounds for MLPs have long been studied in the literature on learning theory; see, e.g., Bartlett (1996); Anthony et al. (1999). More recently, Bartlett et al. (2017) provided a margin bound and corresponding Rademacher analysis that can be adapted to various operator norms at each layer. The proof technique therein was inspired by the early result for two-layer networks in Bartlett (1996). Farrell et al. (2021) established non-asymptotic bounds for MLPs for a general class of nonparametric regression-type loss functions. Schmidt-Hieber (2020) showed that MLPs with a sparsely connected structure achieve the minimax rates of convergence (up to a logarithmic factor) under a general composition assumption on the regression function.
The contribution of the present work is to provide a generalization bound for the novel KAN network structure, which is fundamentally different from the traditional MLPs as demonstrated in Liu et al. (2024c, b). The current analysis uses covering numbers and is closely related to earlier covering number bounds in Anthony et al. (1999) and the recent results in Bartlett et al. (2017) for MLPs. The current proof also utilizes Maurey’s sparsification lemma. However, it is important to highlight some key differences with existing results for MLPs. Firstly, our results can be seen as a non-trivial extension of MLP results to a more flexible network structure specified through a more general choice of activation functions. Secondly, Bartlett et al. (2017) focuses on the margin-based multiclass generalization bound, which has a bounded loss function (more precisely, the ramp loss). In contrast, our results cover a wider range of loss functions (including the ramp loss for classification and a general class of regression-type loss functions), which are not required to be bounded. Thirdly, the results for KANs with a low-rank structure in Section 2.3 appear to be new, and we are not aware of comparable results for MLPs in the recent literature.
The rest of the article is organized as follows: We provide the generalization bounds for KANs equipped with activation functions that are either represented by a set of basis functions or lying in a low-rank space in Sections 2.2 and 2.3, respectively. Section 3 presents numerical results to support the theoretical findings. Section 4 concludes and discusses a few open problems for future research.
@@ 2.2 Risk analysis: activation functions represented by basis functions
As an implication of the above theorem, consider the empirical risk minimization problem in (2), where =$\mathcal{M}=\{f(\cdot)=\bm{\Psi}_{L}\circ\bm{\Psi}_{L-1}\circ\cdots\circ\bm{\Psi}_{1}(\cdot):\bm{\Psi}_{l}\in\mathcal{F}_{l},l=1,2,\dots,L\}$ with $\mathcal{F}_{l}$ specified in Assumption 2. Due to the uniformity over the class $\mathcal{M}$, the risk bound in Theorem 3 holds for any solution to Problem (2) resulting from an optimization scheme (assuming that the resulting solution still belongs to $\mathcal{M}$).
@@
2.3 Risk analysis: activation functions in low-rank RKHS
@@
$R(\widehat{f})-R(f^{*})\leq$ $\frac{6\widetilde{C}^{\prime}(\xi_{0})^{\nu/\tilde{d}}}{n^{(\nu/\tilde{d}+1)/2}(\tilde{d}/\nu-1)^{\nu/\tilde{d}}}(1+\eta^{-1/2})\sqrt{\frac{2(C^{\prime})^{2/s}\log(2/\epsilon)}{n}}$
@@ 3 Numerical Studies
We used both the simulated and real datasets to demonstrate the relationship between the excess loss and the complexity of KANs. We constructed the KANs such that the output of each layer at $\mathbf{0}$ is $\mathbf{0}$, i.e., $\bm{\Psi}_{i}(\mathbf{0})=\mathbf{0}$. Therefore, we used $(\prod_{j=1}^{L}\rho_{j})^{2/3}\sum^{L}_{i=1}(B_{i}c_{i})^{2/3},$ which was proportional to $\tilde{\alpha}$, as the measure of complexity, where the Lipschitz constants $\rho_{j}$s were estimated by their upper bounds provided in Remark 5.
For the simulation, we generated $x_{i}$ from $\sim\text{Unif}(-1,1)$ independently, and let
@@
We run 1000 epochs for each dataset. The results are shown in Figure 2, where we normalize the values of the complexity measures (note that the normalization does not affect the shape of the KAN complexity curve) so that the two curves can be presented in the same plot. The plots in Figure 2 illustrate that our proposed measure of complexity for the KAN networks, $\tilde{\alpha}$, tightly correlates with the excess loss in all the cases. It is surprising to observe that the complexity measure curve closely follows the shape of the excess loss. These results highlight the practical relevance of the generalization bounds derived in Section 2.
@@ 4 Discussions
To conclude, we mention a few future research directions. (i) It would be interesting to derive a lower bound and see how well it matches the upper bound. Also, additional work is needed to determine which norms are well-adapted to KANs as used in practice; (ii) A more careful analysis is needed to show that SGD applied to KAN results in a well-behaved predictor and leads to a refined generalization bound; (iii) Theorem 3 may be improved with a relaxed restriction on the Lipschitz constants of the basis functions $\{g_{ij}^{(l)}\}$; (iv) It was observed that explicit regularization contributes little to the generalization performance of neural networks (Zhang et al., 2021). It is of interest to see if a similar conclusion holds for KANs and whether other types of regularization help to improve the generalization performance.
@@ Appendix A Appendix
A.1 Auxiliary lemmas

### v2

@@ 1.1 Contributions
  • We empirically investigate these bounds when the KANs are trained with stochastic gradient descent (SGD) on simulated and real data sets. The results shed new light on how the performance of KANs is affected by the underlying network structure and the network complexity. In particular, we observe that the complexity measure derived in Section 2 is tightly correlated with the excess loss, demonstrating the complexity measure’s practical relevance. This theoretical-empirical connection suggests a potential strategy for regularization, where the complexity measure could be used as a regularizer in designing the network structure and during the training process, either explicitly or implicitly.
@@ 1.2 Related works
Generalization bounds for MLPs have long been studied in the literature on learning theory; see, e.g., Bartlett (1996); Anthony et al. (1999). More recently, Bartlett et al. (2017) provided a margin bound and corresponding Rademacher analysis that can be adapted to various operator norms at each layer. The proof technique therein was inspired by the early result for two-layer networks in Bartlett (1996). Farrell et al. (2021) established non-asymptotic bounds for MLPs for a general class of nonparametric regression-type loss functions. Schmidt-Hieber (2020) showed that MLPs with a sparsely connected structure achieve the minimax rates of convergence (up to a logarithmic factor) under a general composition assumption on the regression function. The aforementioned work has demonstrated that combining empirical process tools with considerations specific to deep neural networks (DNNs) can yield meaningful results. However, the practical relevance of these results for guiding DNN construction and training requires further investigation. Additionally, alternative theoretical tools have been developed to analyze DNNs, including the neural tangent kernel Jacot et al. (2018) and random matrix theory Mei & Montanari (2022).
The contribution of the present work is to provide a generalization bound for the novel KAN network structure, which is fundamentally different from the traditional MLPs as demonstrated in Liu et al. (2024c, b). The current analysis uses covering numbers and is closely related to earlier covering number bounds in Anthony et al. (1999) and the recent results in Bartlett et al. (2017) for MLPs. The current proof also utilizes Maurey’s sparsification lemma with a suitable adaption to the current setting. However, it is important to highlight some key differences with existing results for MLPs. Firstly, our results can be seen as a non-trivial extension of MLP results to a more flexible network structure specified through a more general choice of activation functions. Secondly, Bartlett et al. (2017) focuses on the margin-based multiclass generalization bound, which has a bounded loss function (more precisely, the ramp loss). In contrast, our results cover a wider range of loss functions (including the ramp loss for classification and a general class of regression-type loss functions), which are not required to be bounded. Thirdly, the results for KANs with a low-rank structure in Section 2.3 appear to be new, and we are not aware of comparable results for MLPs in the recent literature.
The rest of the article is organized as follows: We provide the generalization bounds for KANs equipped with activation functions that are either represented by a set of basis functions or lying in a low-rank space in Sections 2.2 and 2.3, respectively. Section 3 presents numerical results to support the theoretical findings. Section 4 concludes and discusses a few open problems for future research. The appendix contains all the proofs, additional discussions and numerical results.
@@ 2.2 Risk analysis: activation functions represented by basis functions
In Section A.1, we compare this bound with a corresponding bound for MLPs in Bartlett et al. (2017).
@@
As an implication of the above theorem, consider the empirical risk minimization problem in (2), where =$\mathcal{M}=\{f(\cdot)=\bm{\Psi}_{L}\circ\bm{\Psi}_{L-1}\circ\cdots\circ\bm{\Psi}_{1}(\cdot):\bm{\Psi}_{l}\in\mathcal{F}_{l},l=1,2,\dots,L\}$ with $\mathcal{F}_{l}$ specified in Assumption 2. Due to the uniformity over the class $\mathcal{M}$, the risk bound in Theorem 3 holds for any solution to Problem (2) resulting from an optimization scheme (assuming that the resulting solution still belongs to $\mathcal{M}$). In Section A.2, we derive an improved high probability bound when $G(\mathbf{x},y)$ and $B(y)$ both have sub-exponential tails.
@@
2.3 Risk analysis: activation functions in low-rank Sobolev Space
@@
$R(\widehat{f})-R(f^{*})\leq$ $\frac{6\widetilde{C}^{\prime}(\xi_{0})^{\nu/\tilde{d}}}{n^{(\nu/\tilde{d}+1)/2}(\tilde{d}/\nu-1)^{\nu/\tilde{d}}}+(1+\eta^{-1/2})\sqrt{\frac{2(C^{\prime})^{2/s}\log(2/\epsilon)}{n}}$
@@ 3 Numerical Studies
We used both the simulated and real datasets to demonstrate the relationship between the excess loss (defined as the difference between the test loss and training loss) and the complexity of KANs. We constructed the KANs such that the output of each layer at $\mathbf{0}$ is $\mathbf{0}$, i.e., $\bm{\Psi}_{i}(\mathbf{0})=\mathbf{0}$. Therefore, we used $(\prod_{j=1}^{L}\rho_{j})^{2/3}\sum^{L}_{i=1}(B_{i}c_{i})^{2/3},$ which was proportional to $\tilde{\alpha}$, as the measure of complexity, where the Lipschitz constants $\rho_{j}$s were estimated by their upper bounds provided in Remark 5.
For the simulation, we consider two examples in Liu et al. (2024c). Specifically, we generated $x_{i}$ from $\text{Unif}(-1,1)$ independently, and let
@@
We run 1000 epochs for each dataset. The results are shown in Figure 2, where we normalize the values of the complexity measures so that the maximum value of the complexity measure is equal to the last value of the excess loss (see Section A.4 for more details). The plots in Figure 2 illustrate that our proposed measure of complexity for the KAN networks, $\tilde{\alpha}$, tightly correlates with the excess loss in all the cases. It is surprising to observe that the complexity measure curve closely follows the shape of the excess loss. These results highlight the practical relevance of the generalization bounds derived in Section 2. We refer the readers to the appendix for additional numerical results and discussions.
@@ 4 Discussions
To conclude, we mention a few future research directions. (i) It would be interesting to derive a lower bound and see how well it matches the upper bound to understand the tightness of the derived bounds. Also, additional work is needed to determine which norms are well-adapted to KANs as used in practice; (ii) A more careful analysis is needed to show that SGD applied to KAN results in a well-behaved predictor and leads to a refined generalization bound; (iii) Theorem 3 may be improved with a relaxed restriction on the Lipschitz constants of the basis functions $\{g_{ij}^{(l)}\}$; (iv) It was observed that explicit regularization contributes little to the generalization performance of neural networks (Zhang et al., 2021). It is of interest to see if a similar conclusion holds for KANs and whether other types of regularization help to improve the generalization performance. Our theoretical and empirical results suggest a close connection between the derived complexity measure and the generalization error. It is of interest to develop (implicit) regularization techniques to control this complexity measure to achieve improved generalization performance.
@@ Appendix A Appendix
A.1 Comparison with MLPs
We compare the derived bounds with those for MLPs, specifically focusing on the bounds discussed in Bartlett et al. (2017), which address the classification problem with ramp loss. The scenario described in that work closely aligns with the setting of Theorem 2, particularly with $C=0$ and $\max_{i}B(y_{i})$ being finite. Ignoring smaller order terms, the bound for $R(f)-n^{1/2}\sum^{n}_{i=1}\mathcal{L}(f({\mathbf{x}}_{i}),y_{i})$ in Theorem 2 can be expressed as follows
$O\left(\frac{\|\mathbf{X}\|R_{\text{KAN}}}{n}\log(n)\log(2\tilde{d}\tilde{p})+\sqrt{\frac{\log(1/\epsilon)}{n}}\right),$
where
$R_{\text{KAN}}=\left(\prod^{L}_{j=1}\rho_{j}\right)\left(\sum^{L}_{i=1}(B_{i}c_{i})^{2/3}\right)^{3/2}$
measures the KAN complexity. The corresponding bound for MLPs, derived in Theorem 1.1 of Bartlett et al. (2017), is given by
$O\left(\frac{\|\mathbf{X}\|R_{\mathcal{A}}}{\gamma n}\log(n)\log(W)+\sqrt{\frac{\log(1/\epsilon)}{n}}\right),$
where $R_{\mathcal{A}}$ measures model complexity (similar to $R_{\text{KAN}}$ as defined above), $\gamma$ represents the margin in the ramp loss, and $W$ denotes the network width (analogous to $\tilde{d}$ in our case).
When $R$ and $R_{\mathcal{A}}$ are of the same order (denoted by $R\asymp R_{\mathcal{A}}$), both bounds are also of the same order, up to some logarithmic factors. This is expected since $R\asymp R_{\mathcal{A}}$ indicates that both classes exhibit similar levels of complexity. Conversely, if $R=o(R_{\mathcal{A}})$ and $R_{\mathcal{A}}/\sqrt{n}\to+\infty$, then the bound for KANs is of a smaller order than that for MLPs. This scenario arises when the true underlying regression function can be more accurately approximated using KANs. For instance, when the true regression function can be exactly represented by a KAN network, the complexity required for approximating such a function using MLPs is expected to be higher. As discussed in Liu et al. (2024c), KANs may require fewer parameters than MLPs to approximate the same underlying function, as KANs utilize the inherently low-dimensional representation of the true function, unlike MLPs. We refer the readers to Liu et al. (2024c) for examples where KANs can outperform MLPs with the same model complexity, e.g., Figures 3.1-3.3 therein.
A.2 Improved Bounds Under Sub-exponential tails
We derive an improved high probability bound for the generalization errors under the condition that $G(\mathbf{x},y)$ and $B(y)$ both have sub-exponential tails. We impose the following assumption.
Assumption 6.
Given $v\in\mathcal{Y}$, $\mathcal{L}(\cdot,v)$ is Lipschitz in the sense that $|\mathcal{L}(u,v)-\mathcal{L}(u^{\prime},v)|\leq B(v)\|u-u^{\prime}\|_{2}$ for any $u,u^{\prime}\in\mathcal{Y}$ and $B(\cdot):\mathcal{Y}\rightarrow\mathbb{R}_{>0}$. Further suppose $\sup_{f\in\mathcal{M}}|\mathcal{L}(f(\cdot),\cdot)|\leq G(\cdot,\cdot)$. Both $G({\mathbf{x}},y)$ and $B(y)$ have sub-exponential tails, i.e., $\mathbb{E}[\exp(\lambda_{G}G({\mathbf{x}},y))]\leq\exp(\lambda^{2}_{G}\sigma_{G}^{2}/2)$ and $\mathbb{E}[\exp(\lambda_{B}B(y))]\leq\exp(\lambda^{2}_{B}\sigma_{B}^{2}/2)$ for $\lambda_{G},\lambda_{B}$ in a neighborhood of 0.
Theorem 6.
Under Assumptions 1,2 and 6, we have with probability greater than $1-\epsilon-\tau-\eta$,
$R(f)\leq$ $\frac{1}{n}\sum^{n}_{i=1}\mathcal{L}(f(\mathbf{x}_{i}),y_{i})+\frac{144\sqrt{\zeta_{0}}\{\log(C^{\prime}n\log(n)/\sqrt{\zeta_{0}})\vee 1\}}{n}$
$+\frac{C^{\prime}\log(n)\sqrt{\log(2/\epsilon)}}{\sqrt{n}}+\frac{C^{\prime}\log(n)\log(2/\epsilon)}{3n}+C^{\prime}\sqrt{\frac{\log(1/\eta)}{n}},$
for $\epsilon,\tau,\eta>0$ and any $f\in\mathcal{M}$, where $\zeta_{0}=\tilde{\alpha}^{3}\log(2\tilde{d}\tilde{p})(C^{\prime}\log(nC^{\prime}/\tau))^{2}$ and $C^{\prime}$ is a large enough constant.
Note that the dependency $1/\eta$ in the original bound has been improved to $\sqrt{\log(1/\eta)}$ by using Bernstein’s inequality. Additionally, the term $n^{(1-s)/(2s)}$ has been refined to $n^{-1/2}$ due to the exponential tails. In a similar fashion, we have the following result, which strengthens Corollary 1.
Corollary 3.
Let $\widehat{f}\in\mathcal{M}$ satisfy that $\sum^{n}_{i=1}\mathcal{L}(\widehat{f}({\mathbf{x}}_{i}),y_{i})\leq\sum^{n}_{i=1}\mathcal{L}(f^{*}({\mathbf{x}}_{i}),y_{i})$. Suppose Assumptions 1,2 and 6 hold. We have with probability greater than $1-\epsilon-\tau-2\eta$,
$R(\widehat{f})-R(f^{*})\leq$ $\frac{144\sqrt{\zeta_{0}}\{\log(C^{\prime}n\log(n)/\sqrt{\zeta_{0}})\vee 1\}}{n}+\frac{C^{\prime}\log(n)\sqrt{\log(2/\epsilon)}}{\sqrt{n}}$
$+\frac{C^{\prime}\log(n)\log(2/\epsilon)}{3n}+C^{\prime}\sqrt{\frac{\log(1/\eta)}{n}},$
for $\epsilon,\tau,\eta>0$, where $\zeta_{0}=\tilde{\alpha}^{3}\log(2\tilde{d}\tilde{p})(C^{\prime}\log(nC^{\prime}/\tau))^{2}$ and $C^{\prime}$ is a large enough constant.
We remark that the bounds in Theorem 5 and Corollary 2 can also be enhanced under the condition of sub-exponential tails for $G({\mathbf{x}},y)$ and $B(y)$.
A.3 A lower bound on the empirical Rademacher complexity
We derive a lower bound on the empirical Rademacher complexity defined in (7) for the network class specified in Assumption 2. From the second layer, suppose the basis functions include the projection onto the first coordinate, i.e., there exists a linear combination of the basis such that $\sum_{j}\beta_{ij}^{(l)}g_{ij}^{(l)}(\mathbf{x})=x_{1}$ for $\mathbf{x}=(x_{1},\dots,x_{d_{l-1}})$. For the first layer, we set $\psi_{1}^{(1)}(\mathbf{x})=\sum^{p_{1}}_{j=1}\beta_{1j}^{(1)}g_{1j}^{(1)}(\mathbf{x})$ and $\psi_{i}^{(1)}(\mathbf{x})=0$ for all $i\geq 2$. Let $\psi_{1}^{(l)}(\mathbf{x})=x_{1}$ and Let $\psi_{i}^{(l)}(\mathbf{x})=0$ for $i\geq 2$ and $11\leq l\leq L-$ 1. The activation function for the output layer is a real-valued function that maps the input to its first coordinate. Note that the condition $|\psi_{1}^{(1)}(\mathbf{x})-\psi_{1}^{(1)}(\mathbf{x}^{\prime})|\leq\rho\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}$ is implied by $\sum^{p_{1}}_{j=1}|\beta_{1j}^{(1)}|\leq\rho/\max_{j}c_{1j}^{(1)}$ as $|g_{1j}^{(1)}(\mathbf{x})-g_{1j}^{(1)}(\mathbf{x}^{\prime})|\leq c_{1j}^{(1)}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}$. Therefore, $\bm{\Psi}_{L}\circ\cdots\circ\bm{\Psi}_{1}(\mathbf{x})=\sum^{p_{1}}_{j=1}\beta_{1j}^{(1)}g_{1j}^{(1)}(\mathbf{x})$ and the Rademacher complexity of the network class is lower bounded by that of the class
$\mathfrak{G}_{1}=\left\{\psi(\mathbf{x})=\sum^{p_{1}}_{j=1}\beta_{1j}^{(1)}g_{1j}^{(1)}(\mathbf{x}):\left(\sum^{p_{1}}_{j=1}|\beta_{1j}^{(1)}|^{2}\right)^{1/2}\leq\widetilde{B}_{1}\right\}$
for $\widetilde{B}_{1}=\min\{B_{1},\rho/\max_{j}c_{1j}^{(1)}\}/\sqrt{p_{1}}$. Then we have
$\mathcal{R}(\mathfrak{G}_{1}(\mathbf{X}))=$ $\frac{1}{n}\mathbb{E}\left[\sup_{\psi\in\mathfrak{G}_{1}}\sum^{n}_{i=1}e_{i}\psi({\mathbf{x}}_{i})\Big{|}{\mathbf{X}}\right]$
$=$ $\frac{1}{n}\mathbb{E}\left[\sup_{\bm{\beta}:\|\bm{\beta}\|_{2}\leq\widetilde{B}_{1}}\sum^{n}_{i=1}e_{i}\sum^{p_{1}}_{j=1}\beta_{j}g_{1j}^{(1)}(\mathbf{x}_{i})\Big{|}{\mathbf{X}}\right]$
$=$ $\frac{\widetilde{B}_{1}}{n}\mathbb{E}\left[\left\|\sum^{n}_{i=1}e_{i}g_{1j}^{(1)}(\mathbf{x}_{i})\right\|_{2}\Big{|}{\mathbf{X}}\right]\geq\frac{C^{\prime}\widetilde{B}_{1}\{\sum^{n}_{i=1}(g_{1j}^{(1)}(\mathbf{x}_{i}))^{2}\}^{1/2}}{n},$
for some constant $C^{\prime}>0$, where we have used the Khintchine inequality to get the last inequality.
A.4 Additional Numerical Results
In our experiments, the activation functions in the implemented KAN are linear combinations of SiLU function and basis splines, as proposed by Liu et al. (2024c). Specifically, the $\psi_{i,k,j}(x_{j})$ for $i=1,\dots,L,k=1,\dots,d_{i},j=1,\dots,d_{i-1}$ in (1) is
$\psi_{i,k,j}(x)=w^{(b)}_{i,k,j}b(x)+w^{(s)}_{i,k,j}\text{spline}(x),$
where
$b(x)=\text{silu}(x)=x/(1+e^{-x}),\quad\text{spline}(x)=\sum_{i}{c_{i}}B_{i}(x),$
and $B_{i}(x)$ are spline basis. See equations (2.10)-(2.12) of Liu et al. (2024c).
In Figure 2, we display the curves of the KAN complexity as measured by $(\prod^{L}_{j=1}\rho_{j})^{2/3}\sum^{L}_{i=1}(B_{i}c_{i})^{2/3}$ and the excess loss on the same plot. Since the values of these two terms are on different scales, we normalize the complexity so that the maximum value of the complexity measure is equal to the last value of the excess loss. Specifically, let $u=(u_{1},\dots,u_{N})$ be the values of the differences between the test losses and training losses, where $N$ is the number of training epochs, and $v=(v_{1},\dots,v_{N})$ be the model complexity corresponding to different training epoch. Let $v_{\text{max}}=\max\{v_{1},\dots,v_{N}\}$ and $v_{\text{min}}=\min\{v_{1},\dots,v_{N}\}$. The normalized complexity is $v^{\prime}=(v^{\prime}_{1},\dots,v^{\prime}_{N})$ with
$v^{\prime}_{i}=\frac{(v_{i}-v_{\text{min}})u_{N}}{v_{\text{max}}-v_{\text{min}}}.$
We present the curves of the KAN complexity, test loss, and training loss in Figure 3.
Figure 3: The training loss, test loss and complexity (normalized with respect to test loss) of KANs trained with SGD on the four simulated datasets (i–iv) and the MNIST and CIFAR10 datasets. The loss for (iii), (iv), MNIST, and CIFAR10 is the cross-entropy loss, and that for (i) and (ii) is the mean squared error.
The aim of our simulation study is to demonstrate the practical relevance of the generalization bound derived in Section 2. Specifically, we hope to understand the relationship between model complexity, excess risk, and test loss. Our empirical results indicate a strong correlation between excess loss/test loss and model complexity. This finding motivates the need to develop regularization techniques, whether explicit or implicit (such as early stopping or penalization), to control this complexity measure, which can potentially enhance generalization performance.
We use the Setups (i) and (iii) for a simple demonstration, where the underlying regression and classification models are both based on the synthetic function $f_{1}$ defined in Section 3. We apply the dropout technique with a rate of 0.1 to the activation functions in KAN networks (referred to as regularized KAN) with the same shapes as we have used in Section 3 (i.e., [4, 50, 100, 50, 1] for both setups). The results are shown in the top two panels of Figure 4, where we also plot the results for non-regularized KANs, i.e., those results in Figure 2, for comparison. The values of excess loss are depicted in their original scales, while the values of the KAN complexity are normalized with respect to their excess losses. We observe that with the dropout technique, the excess losses decrease significantly from 0.33 to 0.03 for Setup (i) and from 2.41 to 0.06 for Setup (iii). For both the regularized KAN and non-regularized KAN, we observe that the model complexity tightly correlates with the excess loss. As the two complexity curves are plotted with different normalization scales, they cannot be compared directly. Therefore, we plot the ratios of the complexities of the regularized KAN to the non-regularized KAN in the bottom two panels of Figure 4. We can see that the regularized KAN, which induces a lower excess loss than the non-regularized KAN, also has a lower complexity.
Figure 4: (Top) The excess loss and (normalized) complexity of KANs trained with SGD on the simulated datasets (i) and (iii). The loss for (i) is the mean squared error, and that for (iii) is the cross-entropy loss. The green and red curves in the top-right panel overlap. (Bottom) The ratio of the complexity of the regularized KAN to the non-regularized KAN.
A.5 Auxiliary lemmas